# Science Opportunities of Wet Extreme Mass-Ratio Inspirals

### v1

@@ II.3 Calibrating other Methods of Probing Black Hole Mass and Spin
0.2 $1.0\times 10^{6}$ 0.2 20 $1.3\times 10^{42}$ $1.1\times 10^{-14}$
@@
Table 2: Exemplary estimations of the $2-10$ keV X-ray luminosity and flux for SMBH counterparts at redshift $z=0.1$.
@@
Object Mass [×$\times 10^{6}M_{\odot}$] Spin $z$ SNR ($z=0.3$) Alternate Names
@@
UGC 01032 $\sim 1.1$ $0.66^{+0.30}_{-0.54}$ 0.01678 87 56 280 Mrk 359
UGC 12163 $\sim 1.1$ $>0.9$ 0.02468 92 61 290 Ark 564
Swift J2127.4+5654 $\sim 1.5$ $0.6\pm 0.2$ 0.01400 66 38 210
NGC 4253 $1.8^{+1.6}_{-1.4}$ $>0.92$ 0.01293 64 37 200 UGC 07344, Mrk 766
NGC 4051 $1.91\pm 0.78$ $>0.99$ 0.00234 33 19 110 UGC 07030
NGC 1365 $\sim 2$ $>0.97$ 0.00545 29 16 92
1H0707-495 $\sim 2.3$ $>0.94$ 0.04056 19 11 63
MCG-6-30-15 $2.9^{+1.8}_{-1.6}$ $0.91^{+0.06}_{-0.07}$ 0.00749 10 6.8 36
NGC 5506 $\sim 5$ $0.93\pm 0.04$ 0.00608 3.0 2.3 11 Mrk 1376
IRAS13224-3809 $\sim 6.3$ $>0.975$ 0.06579 1.7 1.5 6.6
Ton S180 $\sim 8.1$ $>0.98$ 0.06198 0.9 0.9 3.7
The mass and spin measurements of SMBHs in AGNs have been a long-standing challenge. Standard techniques involving spatially resolved kinematics of stars or the $M_{\bullet}-\sigma_{\star}$ relation [77] generally are not applicable because the strong AGN continuum has seriously diluted any features of stars [78, 79]. GWs provide precise (better than $0.1\%$ [2]) measurements of the mass and even spin of SMBHs, which can be used to calibrate traditional EM wave methods. For this purpose, wet EMRIs that take place in type I AGNs are favored, as traditional measurements based on AGN features all have large uncertainties and thus require improved calibration.
@@
To estimate the fraction of wet EMRIs that could be used for such calibration, the active type I fraction of galaxies was estimated in a two-step process. First, following [80], we adopt the active ($\lambda_{\rm Edd}>10^{-3}$) fraction of $\sim 15-40$ for galaxies with $6.55.5<\log_{10}(M_{\bullet})<$ 6.5, where $\lambda_{\rm Edd}$ is the Eddington ratio. This is broadly consistent with the active fraction of $\sim 24-34$ for the same BH mass range estimated by [81], although a lower Eddington ratio is adopted. The active fraction of [80] is comparable to that of $\sim 10-16$ for a larger BH mass range $10.56.5<\log_{10}(M_{\bullet})<$ 10.5 [82], where the same Eddington ratio is adopted. Furthermore, [83] yielded an active fraction of $\sim 10-22$ for the host galaxy stellar mass range $10.09.0<\log_{10}(M_{\rm\star,~{}host})<$ 10.0. Assuming a $M_{\bullet}-M_{\rm\star,~{}host}$ relation (e.g., Eq. 39 of [80]), this translates to a BH mass range of $6.05.0<\log_{10}(M_{\bullet})<$ 6.0. Second, the active fraction of [80] should be multiplied by the type I fraction of active galaxies. The type I fraction of active galaxies ranges from $\sim 7$ [84] to $\sim 20$ [85, 86]. Therefore, the active type I fraction of galaxies is $\sim 1-8$ [80]. This is comparable to that of $\sim 4$ for a larger BH mass range $10.56.5<\log_{10}(M_{\bullet})<$ 10.5 [82]. Hence, merely a few percent of wet EMRIs could be used to calibrate traditional X-ray measurements of the SMBH spin and optical measurements of the SMBH mass.
@@
For optical counterparts of such wet EMRI, the most commonly used method to estimate $M_{\bullet}$ is the so-called virial mass estimator, which is based on costly and time-consuming reverberation mapping or convenient yet less accurate single-epoch spectra [87, 88]. It assumes that the regions emitting broad emissions, known as broad-line region (BLR) clouds, are entirely governed by the gravitational force of the BH, so that $M_{\bullet}$ can be measured as long as the velocity ($v$) and distance ($r$) of the BLR clouds from the BH are known, i.e. $M_{\bullet}=f{{rv^{2}}\over{G}}$. There is a uniformly assigned coefficient $f$ in the calculation, which is calibrated by the $M_{\bullet}-\sigma_{\star}$ relation, assuming that active and inactive galaxies follow the same relation. Note that the intrinsic scatter of the tightest $M_{\bullet}-\sigma_{\star}$ relation for local massive galaxies is at least $0.3-0.5$ dex  [89, 77, 90], the uncertainty associated with the $f$ used for AGNs could be considerably larger. Moreover, for a given AGN, it suffers from additional uncertainties from the kinematics, geometry, inclination of the clouds, and even the nature of the bulges [91], making the real $f$ hugely different from object to object. Recently, some exciting progress has been made with the advanced NIR interferometry GRAVITY, mounted on the Very Large Telescope Interferometer (VLTI), which has shown promise in probing the BLR structure and thus improving the precision of the $M_{\bullet}$ measurement [92, 93]. Unfortunately, it can only be applied to very few near-infrared luminous AGNs.
@@
For those active type I galaxies with wet EMRI signals, their $2-10$ keV luminosity is estimated as
$L_{\rm 2-10~{}keV}=1.3\times 10^{44}\left(\frac{M_{\bullet}}{10^{6}~{}M_{\odot}}\right)\left(\frac{\lambda_{\rm Edd}}{\kappa_{\rm 2-10~{}keV}}\right)$ (21)
where $M_{\bullet}$ is the BH mass, $\lambda_{\rm Edd}$ the Eddington ratio, $\kappa_{2-10}~{}{\rm keV}$ the $2-10$ keV bolometric correction. The ratio of $\lambda_{\rm Edd}/\kappa_{\rm 2-10~{}keV}$ is in the range of $0.001-0.1$ [97]. Generally speaking, $\kappa_{\rm 2-10~{}keV}\sim 20-70$ for $0.20.1<\lambda_{\rm Edd}<$ 0.2 and $\kappa_{\rm 2-10~{}keV}\sim 70-150$ for $\lambda_{\rm Edd}>0.2$ [98]. The $2-10$ keV flux is obtained following the inverse square law of the luminosity distance, where we assume the standard flat $\Lambda$CDM cosmology with parameters $H_{0}=70~{}{\rm km~{}s^{-1}~{}Mpc^{-1}}$ and $\Omega_{m}=1-\Omega_{\Lambda}=0.3$. Table 2 lists exemplary estimations, mainly for $z=0.1$. The estimated $2-10$ keV X-ray flux is low in general ($\lesssim 10^{-13}~{}{\rm erg~{}s^{-1}~{}cm^{-2}}$) but detectable. For comparisons, the XMM-COSMOS survey can reach the $2-10$ keV flux limit of $\sim 9.3\times 10^{-15}~{}{\rm erg~{}s^{-1}~{}cm^{-2}}$, with an average vignetting-corrected exposure of 40 ks [99]. To estimate the BH spin, merely detection is not enough. Ideally, it is preferable to match the spectral quality of the nearby AGN with the BH spin measurement [96]. For comparisons, the $2-10$ keV flux of Mrk 359 ($z=0.01678$, [100]) listed in Table 3 is $\gtrsim 5\times 10^{-12}~{}{\rm erg~{}s^{-1}~{}cm^{-2}}$. For the X-ray counterparts of wet EMRIs, it is essential to take deep-exposure observations using future X-ray missions (e.g., eXTP [101] and Athena [102]) with significantly larger effective areas than those of current missions.
Table 3 presents a catalog of selected SMBHs with masses below $10^{7}M_{\odot}$, whose masses and spins have been determined using EM observations [103, 96, 104]. The table also includes SNRs for GW detections from LISA, TianQin, and Taiji, assuming a source with the same intrinsic parameters placed at the redshift of $z=0.3$ and with four years of observation. Although EM observations of SMBH spin typically carry significant uncertainties due to the limitations of current methods, GW observations from wet EMRIs provide far greater precision, with uncertainties as low as $10^{-6}\sim 10^{-4}$ for mass and $10^{-6}\sim 10^{-3}$ for spin [2]. This comparison highlights the potential of wet EMRIs to calibrate traditional EM techniques, which rely on indirect methods such as X-ray reflection or virial mass estimators, often affected by large scatter and systematic biases. By integrating EM and GW data, Table 3 illustrates the critical role of wet EMRIs in improving the accuracy and reliability of SMBH mass and spin measurements, especially in type I AGNs.
@@ II.4 Testing Accretion Disk and Jet Models
With the BZ mechanism [105] a jet extracts energy and angular momentum from a spinning SMBH through magnetic fields that thread its event horizon. This process converts the rotational energy of the BH into jet power via the Penrose process [106] to jet energy, so that the direction of the jet should be highly correlated with the spin axis of the BH $\hat{a}$.
In contrast, through the BP mechanism [107] a jet extracts energy from the rotational motion of the accretion disk via magnetic fields that thread the disk. These fields centrifugally accelerate material along magnetic field lines and collimate it into a jet [108]. The resulting jet aligns with the direction of the disk magnetic field, which should be highly correlated with the angular momentum $\hat{L}_{\rm disk}$. Notice that $\hat{L}_{\rm disk}$ may differ from $\hat{a}$ due to disk warping or misalignment, as discussed in Sec. II.2.
Understanding the directional relationships among the accretion disk $\hat{L}_{\rm disk}$, the SMBH spin axis $\hat{a}$, and the jet direction (as illustrated in Fig. 2) is crucial for interpreting the jet-launching process. These relationships not only help disentangle the dominant jet-launching mechanism but also provide insights into the interplay between BH spin, accretion disk dynamics, and magnetic fields.
Figure 1: Posterior distribution of the inclination angle $\iota$, with the dimensionless semi-latus rectum $p_{0}\sim 8M_{\bullet}$ (left panel) and traced back to $p\sim 150M_{\bullet}$ (right panel). The orange vertical lines indicate the mean value along with the $1-\sigma$ uncertainty.
Figure 2: The relationship among the orbital angular momentum of the sBH $\hat{L}$, the SMBH spin direction $\hat{a}$, and the jet angle. The disk orientation $\hat{L}_{\rm disk}$ is inferred by tracing $\hat{L}$ backward from the LISA band ($p_{0}\sim 8M_{\bullet}$) to $p\sim 150M_{\bullet}$, where the sBH resides within the disk.
@@
The spin axis of the SMBH ($\hat{a}$) is the most straightforward to measure of the three vectors, thanks to the exceptional precision of GW observations. GW observations offer precise measurements of the SMBH’s spin, including both its magnitude and direction, with an accuracy ranging from $10^{-6}\sim 10^{-3}$ [2], which is more than three orders of magnitude better than the precision achieved through EM observations (as summarized in Table 3). Moreover, GW signals provide highly accurate measurements of the inclination angle $\iota$, which represents the angle between the SMBH spin axis ($\hat{a}$) and the orbital angular momentum vector ($\hat{L}$), at $p\sim 8M_{\bullet}$, corresponding to the LISA band.
In the warped disk scenario, determining the orientation of the accretion disk ($\hat{L}_{\rm disk}$) is more challenging, as the evolution of sBH depends on a complex interplay of forces acting at different radii. In the outer regions of the disk, the motion of the sBH is primarily governed by disk forces, while closer to the SMBH, GW radiation dominates its evolution. At larger radii, the sBH moves within the plane of the disk, but as it approaches smaller radii, it gradually departs from the disk due to Lense-Thirring precession induced by the SMBH [109].
To reconstruct the orientation of the accretion disk at larger radii, GW measurements of the orbital configuration of sBH in the LISA band (with $p\sim 8M_{\bullet}$) can be used as a starting point to evolve the system backward to $p\sim 150M_{\bullet}$, where the GW lifetime becomes 0.5 Myrs and the disk forces begin to dominate. For simplicity, in this backward evolution, we take a crude approximation that GW damping is the only mechanism that governs the dynamics of the system. At $p\sim 150M_{\bullet}$, the orientation of the disk, influenced by Lense-Thirring precession, can be decomposed into two components: the angle of inclination relative to the SMBH spin axis and the angle of precession, which describes its azimuthal orientation. This approach enables a reconstruction of the orientation of the disk and provides critical insights into its role in the jet-launching process.
To determine the inclination angle of the accretion disk relative to the SMBH’s spin, we first estimate the inclination angle $\iota$, defined as [110]
@@
where $L_{z}$ represents the $z$-component of the orbital angular momentum of the sBH, and $Q$ is the Carter constant. The inclination angle $\iota$ represents the orientation of the orbital angular momentum relative to the spin of SMBH. However, as $\iota$ increases over time due to the radiation reaction [111, 112, 113], it no longer reliably indicates the orientation of the accretion disk. To estimate the disk’s orientation, the orbital trajectory must be evolved backward to regions far from the SMBH, where $\iota$ can, in principle, reflect the orientation of the accretion disk.
The measurement uncertainty of $\iota$ as the EMRI system enters the LISA band ($p_{0}\sim 8M_{\bullet}$) is estimated by performing Fisher analysis with the Augmented Analytic Kludge (AAK) waveform model [46, 47]. Fisher analysis is a well-established statistical method for quantifying the precision with which GW observatories, such as LISA, can extract source parameters from observed signals. The Fisher information matrix, $\Gamma_{ij}$, is defined as:
where $h$ is the signal, $\lambda_{i}$ are the source parameters, and the inner product is defined by Eqn. 7. The covariance matrix, $\Sigma_{ij}=(\Gamma^{-1})_{ij}$, is obtained by inverting the Fisher matrix, with the diagonal elements, $\sigma_{\lambda_{i}}^{2}=\Sigma_{ii}$, quantify the variances of the parameters.
The posterior samples obtained from this analysis are propagated backward by evolving the orbital trajectory to $p\sim 150M_{\bullet}$, where the orbital angular momentum of the sBH provides an estimate of the accretion disk’s orientation and, consequently, the angle relative to the SMBH’s spin. As an example, we use a simulated EMRI system with the following key parameters:
@@
As shown in Fig. 1, the $1-\sigma$ uncertainty for the inclination angle is approximately $\sim 0.05^{\circ}$. This angle corresponds to the inclination of the warped accretion disk at the decoupling radius ($p\sim 150M_{\bullet}$), where the influence of disk torques begins to dominate over GW radiation. This result demonstrates that GWs provide a highly precise measurement of the disk’s inclination at the critical decoupling radius, as discussed in Sec. II.2.
On the other hand, to estimate the uncertainty in the precession angle at the decoupling radius, we perform similar analysis to propagate samples of orbital configurations measured at $p_{0}\sim 8M_{\bullet}$ (within the LISA band) backward to a semi-latus rectum of $p\sim 150M_{\bullet}$, where the orbital direction is measured. It is much larger than $R_{\rm warp}$ (see Eq. 8) so $p\sim 150M_{\bullet}$ already corresponds to the outer-disk region. The uncertainties in the precession angle grow approximately linearly with increasing $p$, reaching a value of around $10^{\circ}$ at $p\sim 150M_{\bullet}$ as shown in Fig. 3. This level of uncertainty makes it challenging to resolve the azimuthal orientation of the warped disk.
In addition to GW-derived measurements of the SMBH spin and the inclination angle, EM observations can estimate the jet direction if the host AGN of the EMRI is resolvable. For AGNs with jet viewing angles smaller than $\sim 10^{\circ}$, radio observations, such as those conducted by the MOJAVE program, can determine the jet angle with an accuracy of approximately $1^{\circ}$ by analyzing jet kinematics [114, 115, 116, 117, 118, 119, 120, 121]. Furthermore, for gamma-ray bright AGNs, observations from Fermi-LAT typically measure larger jet viewing angles and greater accuracy on average compared to non-LAT-detected AGNs [115, 118].
The interplay between the jet direction, the SMBH spin axis, and the orientation of the accretion disk will likely provide critical information to understand the mechanisms of accretion and jet launch in AGN. As shown in Fig. 2, GW signals enable precise measurements of the SMBH spin axis and the inclination angle $\iota$, and hence the orientation of the accretion disk relative to the SMBH spin. EM observations, on the other hand, can determine the jet direction. Although this allows for an estimation of the relative angle between the jet and the SMBH spin axis, the angle between the jet and the accretion disk orientation remains inaccessible because of the large uncertainty in the precession angle. This indirect relationship between the accretion disk orientation and the jet direction limits our ability to fully understand the correlations between these vectors. In addition, to statistically establish the correlations, it will likely require studying a population of wet EMRIs (those with significant disk interactions) with well-resolved EM counterparts.

### v2

@@ II.3 Calibrating other Methods of Probing Black Hole Mass and Spin
0.2 $1.0\times 10^{6}$ 0.2 20 $3\times 10^{42} 1.1$$1\times 10^{-14}$
@@
Table 2: Exemplary estimations of the 2102-10 keV X-ray luminosity and flux for SMBH counterparts at redshift $z=0.1$.
@@
Object Mass $\times 10^{6}M_{\odot}$] Spin $z$ SNR ($z=0.3$) Alternate Names
@@
UGC 01032 $\sim 1.1$ $66^{+0.30}_{-0.54} 0.0$1678 87 56 280 Mrk 359
UGC 12163 $\sim 1.1$ >0.9>0.9 0.02468 92 61 290 Ark 564
Swift J2127.4+5654 $\sim 1.5$ $6\pm 0.2 0.0$1400 66 38 210
NGC 4253 $1.8^{+1.6}_{-1.4}$ >0.92>0.92 0.01293 64 37 200 UGC 07344, Mrk 766
NGC 4051 $1.91\pm 0.78$ >0.99>0.99 0.00234 33 19 110 UGC 07030
NGC 1365 $\sim 2$ >0.97>0.97 0.00545 29 16 92
1H0707-495 $\sim 2.3$ >0.94>0.94 0.04056 19 11 63
MCG-6-30-15 $2.9^{+1.8}_{-1.6}$ $91^{+0.06}_{-0.07} 0.0$0749 10 6.8 36
NGC 5506 $\sim 5$ $93\pm 0.04 0.0$0608 3.0 2.3 11 Mrk 1376
IRAS13224-3809 $\sim 6.3$ >0.975>0.975 0.06579 1.7 1.5 6.6
Ton S180 $\sim 8.1$ >0.98>0.98 0.06198 0.9 0.9 3.7
EMRI GWs provide precise (better than $0.1\%$ [2, 68]) measurements of the mass and spin of SMBHs, which can be used to calibrate traditional EM methods if the host AGN can be identified. As discussed in the previous subsections and summarized in Table 1, the host AGN identification are possible for a few to a few tens of EMRI host AGNs.
Generally speaking, traditional EM wave methods perform better for type I AGN than type II AGN because the latter is obscured [96, 97, 98]. While wet EMRIs can occur in type I galaxies with $4.0<\log_{10}(M_{\bullet})<7.0$, only a small fraction of galaxies in this mass range are active. Limited by available observational studies, we further limit the mass range to $5.5<\log_{10}(M_{\bullet})<7.0$.
@@
The active type I fraction of galaxies with $5.5<\log_{10}(M_{\bullet})<7.0$ was estimated in a two-step process. First, following [99], we adopt the active ($\lambda_{\rm Edd}>10^{-3}$) fraction of $\sim 15-40$ % for galaxies with $5.5<\log_{10}(M_{\bullet})<6.5$, where $\lambda_{\rm Edd}$ is the Eddington ratio. This is broadly consistent with the active fraction of $\sim 24-34$ % for the same BH mass range estimated by [100], although a lower Eddington ratio is adopted. For a larger BH mass range $6.5<\log_{10}(M_{\bullet})<10.5$, the active ($\lambda_{\rm Edd}>10^{-3}$) fraction is $\sim 10-16$ % [101]. Furthermore, [102] yielded an active fraction of $\sim 10-22$ % for the host galaxy stellar mass range $9.0<\log_{10}(M_{\rm\star,~host})<10.0$. Assuming a $M_{\bullet}-M_{\rm\star,~host}$ relation (e.g., Eqn. 39 of [99]), this translates to a BH mass range of $5.0<\log_{10}(M_{\bullet})<6.0$.
Second, the active fraction of [99] should be multiplied by the type I fraction of active galaxies. The type I fraction of active galaxies ranges from $\sim 7$ % [103] to $\sim 20$ % [104, 105]. Therefore, the active type I fraction of galaxies with $5.5<\log_{10}(M_{\bullet})<6.5$ is $\sim 1-8$ % [99]. This is comparable to that of $\sim 4$ % for a larger BH mass range $6.5<\log_{10}(M_{\bullet})<10.5$ [101]. Hence, a few percent of galaxies with $5.5<\log_{10}(M_{\bullet})<7.0$ might be used to calibrate traditional X-ray measurements of the SMBH spin and optical measurements of the SMBH mass using precise GW spin and mass measurement.
@@
For optical counterparts of such wet EMRI, the most commonly used method to estimate $M_{\bullet}$ is the so-called virial mass estimator, which is based on costly and time-consuming reverberation mapping or convenient yet less accurate single-epoch spectra [106, 107]. It assumes that the regions emitting broad emissions, known as broad-line region (BLR) clouds, are entirely governed by the gravitational force of the BH, so that $M_{\bullet}$ can be measured as long as the velocity ($v$) and distance ($r$) of the BLR clouds from the BH are known, i.e. $M_{\bullet}=f{{rv^{2}}\over{G}}$. There is a uniformly assigned coefficient $f$ in the calculation, which is calibrated by the $M_{\bullet}-\sigma_{\star}$ relation, assuming that active and inactive galaxies follow the same relation. Note that the intrinsic scatter of the tightest $M_{\bullet}-\sigma_{\star}$ relation for local massive galaxies is at least 0.30.50.3-0.5 dex  [108, 109, 110], the uncertainty associated with the $f$ used for AGNs could be considerably larger. Moreover, for a given AGN, it suffers from additional uncertainties from the kinematics, geometry, inclination of the clouds, and even the nature of the bulges [111], making the real $f$ hugely different from object to object. Recently, some exciting progress has been made with the advanced NIR interferometry GRAVITY, mounted on the Very Large Telescope Interferometer (VLTI), which has shown promise in probing the BLR structure and thus improving the precision of the $M_{\bullet}$ measurement [112, 113]. Unfortunately, it can only be applied to very few near-infrared luminous AGNs.
@@
For those active type I galaxies with wet EMRI signals, their 2102-10 keV luminosity is estimated as
$L_{\rm 2-10~keV}=1.3\times 10^{44}{\rm~erg~s^{-1}}\left(\frac{M_{\bullet}}{10^{6}~M_{\odot}}\right)\left(\frac{\lambda_{\rm Edd}}{\kappa_{\rm 2-10~keV}}\right)$ (25)
[…]
Table 3 presents a catalog of selected SMBHs with masses below $10^{7}M_{\odot}$, whose masses and spins have been determined using EM observations [123, 116, 124]. The table also includes SNRs for GW detections from LISA, TianQin, and Taiji, assuming a source with the same intrinsic parameters placed at the redshift of $z=0.3$ and with four years of observation. Although EM observations of SMBH spin typically carry significant uncertainties due to the limitations of current methods, GW observations from wet EMRIs provide far greater precision, with uncertainties as low as $10^{-6}\sim 10^{-4}$ for mass and $10^{-6}\sim 10^{-3}$ for spin [2, 68]. This comparison highlights the potential of wet EMRIs to calibrate traditional EM techniques, which rely on indirect methods such as X-ray reflection or virial mass estimators, often affected by large scatter and systematic biases. By integrating EM and GW data, Table 3 illustrates the critical role of wet EMRIs in improving the accuracy and reliability of SMBH mass and spin measurements, especially in type I AGNs.
@@ II.4 Testing Accretion Disk and Jet Models
With the BZ mechanism [125], a jet extracts energy and angular momentum from a spinning SMBH through magnetic fields that thread its event horizon. This process converts the rotational energy of the BH into jet power via the Penrose process [126] to jet energy, so that the direction of the jet should be highly correlated with the spin axis of the BH $\hat{a}$.
In contrast, through the BP mechanism [127], a jet extracts energy from the rotational motion of the accretion disk via magnetic fields that thread the disk. These fields centrifugally accelerate material along magnetic field lines and collimate it into a jet [128]. The resulting jet aligns with the direction of the disk magnetic field, which should be highly correlated with the angular momentum $\hat{L}_{\rm disk}$. Notice that $\hat{L}_{\rm disk}$ may differ from $\hat{a}$ due to disk warping or misalignment, as discussed in Sec. II.2.
Understanding the geometric relationships among the SMBH spin axis $\hat{a}$, the accretion disk $\hat{L}_{\rm disk}$ and the jet direction (as illustrated in FIG. 3) is crucial for interpreting the physical processes behind the jet-launching. Disentangling these directional vectors allows us to constrain the dominant jet-launching mechanism and to better understand the interplay between SMBH spin, disk dynamics, and magnetic fields.
Figure 2: Posterior distribution of the inclination angle $\iota$, shown for the dimensionless semi-latus rectum at $p_{0}\sim 8M_{\bullet}$ (left panel), and after tracing the evolution back to $p\sim 150M_{\bullet}$ due to GW radiation, and further to $p=410M_{\bullet}$ using Eqn. 14 with $\beta$-disk model (right panel). The orange vertical lines indicate the mean value along with the $1-\sigma$ uncertainty.
Figure 3: The relationship among the orbital angular momentum of the sBH $\hat{L}$, the SMBH spin direction $\hat{a}$, and the jet angle. The disk orientation $\hat{L}_{\rm disk}$ is inferred by tracing $\hat{L}$ backward from the LISA band ($p_{0}\sim 8M_{\bullet}$) to $p\sim 410M_{\bullet}$ (decoupling radius), where the sBH resides within the disk.
@@
Among the three vectors, the SMBH spin axis $\hat{a}$ is the most straightforward to measure using GW observations. The LISA mission is expected to constrain the direction of the SMBH spin with an angular uncertainty in the range of approximately $1-10\,\rm deg^{2}$ for $z<0.25$ [71]. Additionally, GW signals allow accurate extraction of the inclination angle $\iota$, defined as the angle between the SMBH spin axis ($\hat{a}$) and the orbital angular momentum vector ($\hat{L}$). This measurement is typically made at a semi-latus rectum of $p\sim 8M_{\bullet}$, corresponding to the region where LISA is most sensitive.
In warped disk scenarios, inferring the orientation of the accretion disk ($\hat{L}_{\rm disk}$) is more challenging, as the evolution of sBH depends on a complex interplay of forces acting at different radii (as illustrated in FIG. 1). In the outer regions of the disk, the motion of the sBH is primarily governed by disk forces, while closer to the SMBH, GW radiation dominates its evolution. At larger radii, the sBH orbits within the disk plane, but as it migrates inward, it gradually departs from the disk due to Lense-Thirring precession induced by the SMBH [129].
To estimate the disk orientation at larger radii, we perform a backward evolution of the orbital trajectory—driven solely by GW damping and neglecting other dynamical interactions for simplicity—from the LISA band ($p\sim 8M_{\bullet}$) to $p\sim 150M_{\bullet}$. We then evolve the system further to $p=410M_{\bullet}$ using Eqn. 14 with the $\beta$-disk model. Under this assumption, the orbital angular momentum vector of the sBH, reconstructed at $p\sim 410M_{\bullet}$, serves as a proxy for the accretion disk orientation at that radius. This radius is significantly larger than the warp radius $R_{\rm warp}$ (see Eqn. 11) and corresponds to the outer region of the accretion disk, where disk forces are expected to dominate over relativistic effects, as discussed in Sec. II.2.
At this outer location, the disk orientation can be described by two angles measured relative to the SMBH spin axis: the inclination angle, which captures the tilt of the disk, and the precession angle, which describes its azimuthal orientation. Together, these angles define the full 3D geometry of the disk and provide essential input for understanding the jet-launching mechanism in the context of disk–black hole interactions.
The inclination angle $\iota$ of the sBH’s orbit, defined as [130]
@@
where $L_{z}$ represents the $z$-component of the orbital angular momentum of the sBH, and $Q$ is the Carter constant. The inclination angle $\iota$ represents the orientation of the orbital angular momentum relative to the spin of SMBH. However, as $\iota$ increases over time due to the radiation reaction [131, 132, 133], it no longer reliably indicates the orientation of the accretion disk. Thus, to estimate the disk’s orientation, the orbital trajectory must be evolved backward to regions far from the SMBH, where $\iota$ can, in principle, reflect the orientation of the accretion disk.
The measurement uncertainty of $\iota$ as the EMRI system enters the LISA band ($p_{0}\sim 8M_{\bullet}$) is estimated by performing Fisher analysis with the AAK waveform model [56, 59]. The posterior samples obtained from this analysis are propagated backward by evolving the orbital trajectory first to $p\sim 150M_{\bullet}$, and then to $p\sim 410M_{\bullet}$. At this larger radius, the orbital angular momentum of the sBH serves as a proxy for the orientation of the accretion disk. As an illustrative example, we consider a simulated EMRI system with the following key parameters:
@@
We find that the $1\sigma$ uncertainty in $\iota$ at $p\sim 410M_{\bullet}$ is approximately $\sim 0.05^{\circ}$ (see FIG. 2). This result demonstrates that GW observations can provide a highly precise estimate of the disk inclination.
To estimate the uncertainty in the precession angle, we similarly propagate samples of orbital configurations measured at $p_{0}\sim 8M_{\bullet}$ (within the LISA band) backward to a semi-latus rectum of $p\sim 150M_{\bullet}$. Ideally, the backward propagation should extend to $p\sim 400M_{\bullet}$ as illustrated in FIG. 1. However, for $p>150M_{\bullet}$, the disk begins to dominate the orbital evolution. Beyond this point, the uncertainty no longer improves significantly due to the increasing influence of disk-induced torques. At these outer radii, the precession angle reflects the azimuthal orientation of the disk. However, unlike the inclination, the uncertainty in the precession angle grows approximately linearly with distance, reaching a value of around $10^{\circ}$ already at $p\sim 150M_{\bullet}$, as shown in FIG. 4. This level of uncertainty makes it challenging to resolve the azimuthal orientation of the warped disk.
In addition to GW-derived measurements of the SMBH spin and the inclination angle, EM observations can estimate the jet direction if the host AGN of the EMRI is resolvable. For AGNs with small jet viewing angles ($\sim 10^{\circ}$), radio observations, such as those conducted by the MOJAVE program, can determine the jet angle with an accuracy of approximately $1^{\circ}$ by analyzing jet kinematics [134, 135, 136, 137, 138, 139, 140, 141]. Furthermore, for gamma-ray bright AGNs, observations from Fermi-LAT typically measure larger jet viewing angles and greater accuracy on average compared to non-LAT-detected AGNs [135, 138].
The ability to measure the directions of the SMBH spin, accretion disk, and jet opens a promising avenue for testing jet-launching models. As illustrated in FIG. 3, GW observations provide precise measurements of the spin axis and the disk inclination angle, while EM observations reveal the jet orientation. However, due to the uncertainty in the precession angle, the relative angle between the jet direction and the accretion disk remains inaccessible in individual events. As a result, while we can constrain the angle between the jet and the SMBH spin, which is most relevant for testing the BZ mechanism, we cannot yet directly compare the jet to the disk orientation, limiting our ability to test the BP mechanism on a case-by-case basis. In addition, to statistically establish the correlations, it will likely require studying a population of wet EMRIs (those with significant disk interactions) with well-resolved EM counterparts.